OpenStudy (anonymous):

Here Comes another set :) The length of each side of cube is 5 cms, has been painted red,blue and green on pairs of opposite faces. It is then cut into cubical boxes of each side 1cm. Now I want you to answer these below Questions regarding above Question (:P). 1. How many cubes have one face painted red and once face painted blue and other faces may be painted or unpainted? 2. How many cubes are there which doesn't have red or blue paint on any sides of them?

OpenStudy (anonymous):

Come on guys!!

OpenStudy (anonymous):

w8 a min :P

OpenStudy (anonymous):

okay :P

OpenStudy (anonymous):

@uri then bolo?

OpenStudy (anonymous):

1. 10 2. 125 - 50 -40 = 35 ?

OpenStudy (anonymous):

no @mhmdrz91 im sorry :(

OpenStudy (anonymous):

2. 125 - 50 - 30 =45 ?

OpenStudy (anonymous):

so if my answer is incorrect, it's from my english! and i can not understand question ;)

OpenStudy (anonymous):

second one is right!

OpenStudy (anonymous):

1. 40 ?

OpenStudy (anonymous):

20

OpenStudy (anonymous):

the answer of question 1 is 20? yes 20! why i calculating 4*5 = 40? :D i should sleeping i think :D

OpenStudy (anonymous):

Haha :D